OpenStudy (anonymous):

Solve for x in the following equation 2x^2-24x+54=0 A.) x=-9;-3 B.)x=9;-3 C.)x=9;3 D.)x=-9;3

OpenStudy (anonymous):

divide by 2. you will get x^2 - 12x+27 = 0 notice that 12 = 9 +3 and notice that 9*3 = 27

OpenStudy (anonymous):

so its B

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

thenn it has to be A

OpenStudy (anonymous):

why dont you substitute the values of 9,3,-9,-3 in your equation and see which of them solve it?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

okay well i pluged in C and thats the right answer

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

yaaayyy i got it

OpenStudy (anonymous):

:)